# Believing is Seeing: Unobserved Object Detection using Generative Models

### v2

@@ 3 Unobserved Object Detection
The task of unobserved object detection, as shown in Fig. 1, is to detect objects that are in the scene, but are not captured within the camera frustum. The approaches considered in this paper address this task by predicting a conditional distribution $\mathcal{D}_{\mathcal{I}}$ over a bounded space and set of semantic labels given a single RGB image $\mathcal{I}$, referred to as a spatio-semantic distribution (SSD). This facilitates the inference of the probable locations of unobserved objects outside the visible region captured by the image and its associated camera frustum, using contextual cues from the image. The spatial domain $\mathcal{X}$ defines the support of this distribution. The SSD is an estimate of the ground-truth conditional distribution $\Gamma_{\mathcal{I}}$. While this distribution is not observable, we have access to additional posed images from a realization of this conditional distribution for the purposes of evaluation. For each object label $o$, we denote the associated SSD as $\mathcal{D}_{\mathcal{I}o}$. This maps each $x\in\mathcal{X}$ to a likelihood in $[0,1]$. That is, the likelihood of detecting object $o$ at location $x$. For example, it assigns a likelihood to the presence of a chair one meter in front of and two meters to the right of the camera, which may not be visible in the input image. Note that we consider all distributions to be normalized. We consider distributions over two discretized spatial domains $\mathcal{X}$: (1) a 2D domain $\mathbb{I}$, discretized as a pixel grid, corresponding to an expansion of the input image; and (2) a 3D domain $\mathbb{V}$, discretized as a voxel grid, extending beyond the input camera’s frustum. The 2D domain $\mathbb{I}$ symmetrically extends the $H\times W$ pixel grid of the input image to $H\times W^{\prime}$, where $W^{\prime}>W$. The 3D domain $\mathbb{V}$ expands the frustum to an expanded voxel grid aligned with the input camera pose.
@@ 5.1 Metrics for Unobserved Object Detection
for input image $\mathcal{I}$ and object $o$. The normalization factor $\log|\mathcal{X}|$ makes the metric independent of the grid resolution and ensures that a uniform distribution evaluates to unit entropy. This asymmetric measure penalizes the prediction of low object likelihood where the ground-truth indicates the presence of an object. However, as required, it does not penalize the prediction of high object likelihoods where the ground-truth has not seen an object.
For the remaining metrics, we consider high-likelihood detections above a threshold $\tau$. The thresholded ground truth is denoted by $\mathcal{G}_{\mathcal{I}o}^{\tau}$ for all $x\in\mathcal{X}$ where $\mathcal{G}_{\mathcal{I}o}(x)>\tau$. Let the detection label $y_{\mathcal{I}o}$ be $1$ if $\mathcal{G}_{\mathcal{I}o}^{\tau}$ is non-empty and 00 otherwise. The thresholded prediction $\mathcal{D}_{\mathcal{I}o}^{\tau}$ and label $\hat{y}_{\mathcal{I}o}$ are correspondingly defined over $\mathcal{D}_{\mathcal{I}o}$.
@@ 5.2 Experiment Setup
All models were evaluated on 10 indoor scenes from the RealEstate10k test set [87] and the NYU Depth V2 dataset [46], automatically selected based on object detection frequency. After filtering, each scene has an average of 190 images in RealEstate10k ($720\times 1280$, resized to $360\times 640$ for $\mathbb{I}$, and center-cropped to $360\times 360$ for $\mathcal{I}$) and 1359 images in NYU Depth V2 ($480\times 640$, center-cropped to $360\times 640$ for $\mathbb{I}$ and center-cropped to $360\times 360$ for $\mathcal{I}$). One randomly selected frame per scene was used as the input image $\mathcal{I}$ for evaluation. COLMAP [62] and GLOMAP [49] were used for camera pose estimation (for NYU Depth V2) and for reconstruction to obtain depth maps. Object annotations for each expanded $360\times 640$ image $\mathbb{I}$ were generated using YOLOv8x [54, 27], pre-trained on the MS-COCO dataset [35], retaining those with confidence greater than $\tau_{\text{conf}}$. The 2D ground-truth spatial distributions $\mathcal{G}_{\mathcal{I}o}^{\text{2D}}$ were computed by a softmax normalization of the detected object confidence logits. The 3D ground-truth spatial distributions $\mathcal{G}_{\mathcal{I}o}^{\text{3D}}$ were computed by back-projecting the object confidence logits using the 3D reconstructions, voxelizing with average pooling, then applying softmax normalization. The voxel grid has resolution $20\times 20\times 20$ with grid size $1.0$. The 2.5D ground truth spatial distributions $\mathcal{G}_{\mathcal{I}o}^{\text{2.5D}}$ were computed were computed similarly, by back-projecting the confidence logits within the input camera frustum and using a voxel grid of resolution $10\times 10\times 10$, with a grid size of $1.0$. Note that we only retain non-empty ground-truth distributions in the dataset: if a scene $s$ does not contain an object $o$, the pair $(s,o)$ is not part of the dataset. See Appendices A and B for details on automatic scene preprocessing, filtering, selection, and reconstruction.
@@
The hyperparameters are set as follows: the classification threshold $\tau=1.4/|\mathcal{X}|$, the minimum detection confidence $\tau_{\text{conf}}=0.1$, and the number of subdomains $\mathcal{X}^{i}$ for region-wise accuracy $\ell=3$, corresponding to the $360\times 360$ input image $\mathcal{I}$, the $360\times 140$ domain to the left of the image and the $360\times 140$ domain to the right of the image. For DFM [73] inference, we use autoregressive sampling with 3 timesteps per frame, generating 50 intermediate frames with default settings: a temperature of $0.85$ and a guidance scale of $2.0$. Generating each sample took $25$ minutes per target pose per sample on average; that is, 48 days to generate all samples for each dataset. For Stable Diffusion sampling [52], the Runway SDXL inpainting model was employed via the Huggingface API. Input images were padded by 140 pixels on the left and right sides for masked inpainting. The temperature was set to $0.25$, the prompt guidance scales were drawn from $\{0.25,0.5,0.75,1.0\}$, the seeds from $\{0,42,123,999,2021\}$, and the prompts from a set of $10$, yielding $4\times 5\times 10=200$ samples per scene. For the VLMs, we used ChatGPT-4o [47], Claude-Sonnet-3.5 [3], and Gemini 1.5 Ultra [20] via their official APIs, and LLaVa-v1.6-34b [37, 38] via the Replicate API. The temperature was set to 0.25, the top-p value was set to 1.0 where applicable, and the order of the image–question pairs was randomly shuffled. The VLMs were queried 100 times per object, region, and scene, resulting in 6000 calls each ($20\,\text{images}\times 30\,\text{questions}\times 10\,\text{shuffles}$). Data acquisition and post-processing through these APIs took $\sim 8$h. Complete details are in Appendices C and D. All experiments were run on a single NVIDIA RTX A6000 GPU. Our code, dataset and evaluation protocols will be publicly available for reproducibility. For closed-source VLM APIs, future updates or access restrictions could affect reproducibility.
@@ 5.3 Results
Uniform 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1$ 0.553  $\pm$0.220 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1$ 0.558  $\pm$0.214
@@ 2.5D Task 3D Task
Uniform 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1$ 1.000  $\pm$0.000 1.000  $\pm$0.000 0.000  $\pm$0.000 $1$ 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1$ 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1$
@@
Qualitative results are shown in Fig. 5(d), with the $360\times 360$ center-cropped input image $\mathcal{I}$ shown within the dashed lines in the first column, a top-down view of the DFM-predicted 3D distribution in the second column, the SDXL-predicted 2D distribution in the third column, and the ChatGPT-predicted 2D distribution in the last column. We observe that all methods infer relatively sensible distributions. In particular, the DFM-based model usually predicts high likelihoods near the ground-truth object’s 3D location, including when the object is occluded or out-of-frame, which is a significantly harder task than the 2D prediction task; see, for example, the final row. More qualitative results can be found in Appendix E. Failure cases are shown in Fig. 6, where all models perform poorly and an object category (door) that elicits many false negatives in the object detector. We also observe that the diffusion-based outpainting approach has its own modes of failure, also shown in Fig. 6, where the results are not informative in a scene if the model is not prompted with the label of the object to be detected.
@@ 6 Conclusion
This work has several significant limitations. In particular, the inference time of generative models, especially 3D diffusion models, is currently very large. Since our approaches require many samples to approximate the conditional distribution, this is a significant bottleneck and stymies their use in real-time applications like in robotics. Another limitation is the need for object prompts in outpainting-based methods, since this is likely to bias the obtained empirical conditional distribution. Also, VLM-based evaluations may not be fully reliable, since it is possible these models could have seen and memorized the test images during training. Since the task requires that the model not over-fit to the ground truth realization—which is an instance of the true conditional distribution—there are inherent challenges in measuring performance that reflects accurate modeling of the underlying true distribution that is expected to be unavailable in real-world settings.
While the work presents a promising initial step towards this significant task, it also sets up some open questions. For example, while all of the detection pipelines presented in this work use pre-trained generative models, it is likely that training for this task directly would elicit more accurate predictions. Moreover, the 2D and 3D diffusion models are repurposed from their original task of photorealistic generation, but it is unclear whether high-fidelity generation is desirable for unobserved object detection, given trade-offs between consistency and diversity of conditional generation [4]. One might expect it to be simpler to train a model to directly generate samples from the conditional spatio-semantic distribution at lower resolutions.
Future work could investigate task performance on controlled synthetic datasets where a true conditional distribution is accessible [13, 16]. Richer categories of spatio-semantic features like affordances [24] can prove promising extensions. Additionally, solving unobserved object detection can open up applications to a large category of downstream tasks such as reactive robot control [48], visual search [5], planning [32, 28], navigation [64] and scene exploration [25, 18]. While some applications might be amenable to coarse-grained spatial information, others will require finer spatio-semantic reasoning in 2D and 3D, motivating the current work as a promising step towards developing the machine vision capability of unobserved object detection.
@@ Appendix A Object, Scene and Parameter Selection Methodology
To systematically select diverse scenes, we sampled frames from videos in the RealEstate10k test dataset, which contains world-to-camera poses and pinhole camera intrinsics [87], built using the ORBSLAM pipeline [43, 42]. Due to memory limitations, we downloaded 250 random scenes for analysis. RealEstate10k contains indoor and outdoor scenes but lacks metadata or labels, thus to interpret and filter each scene, our objective was to select representative viewpoints across each scene. We selected these frames, for every scene, at approximately geodesically equidistant intervals in $\mathrm{SE}(3)$, between the two farthest camera-to-world poses for each scene, by computing the geodesic distance in $\mathrm{SE}(3)$ [10]. Using these selected images, per scene, we used the ChatGPT-4o API [47] to classify scenes as indoor or outdoor based on these 10 images per scene. For the indoor scenes, we further utilized the API to categorize them into four types: kitchen, study, bedroom, and living room. This automated labeling ensured that our subsequent analysis remained focused and consistent across the relevant indoor scenes.
@@
We deployed an end-to-end automated object selection pipeline on the randomly acquired scenes, logging the occurrence frequency of detected objects across the subsets of both datasets. From the combined scenes of NYU Depth V2 and RealEstate10k scenes, we selected the 20 most frequently detected object classes among the 80 classes in the COCO dataset [35]. For our analysis, we focused on 10 random object categories from these 20 classes that were present in at least one of the chosen scenes in the combined choices of scenes from both datasets. The COCO object categories analyzed include: refrigerator, TV, bed, chair, sink, oven, book, laptop, couch, and door. We used Yolov8x [54, 27] for its strong adaptability for images of various resolutions and superior generalizability.
@@ Appendix C Additional implementation details
The experiments were conducted on a single machine equipped with an NVIDIA A6000 GPU, an AMD Ryzen Threadripper Pro 5995WX CPU, and 64GB of DDR5 RAM, running Ubuntu 22.04 LTS. From sampling to the final metrics summary, our end-to-end pipeline takes up to 38 minutes utilizing multiprocessing. We will release a comprehensive list of environmental dependencies and configurations for the ease of replication.
@@
For the 3D study, we utilized the aggregated point cloud from the samples from the corresponding poses, and correspondingly all valid 3D points in the ground-truth reconstruction, including occluded parts of the scene, voxelized into a grid of size $20\times 20\times 20$, such that the camera position is at the center of the grid. For the SDXL-based study, taking a similar approach as Chen et al. [12] to lift our representation to 2.5D, we generated metric depths up to the far plane of $10$ units using DepthAnythingv2 [80, 81], selected for its superior cross-domain generalizability and compatibility with our known camera model. This approach provided depths within the frustum, ensuring scale consistency across the analyses, making sure that the camera centers are aligned to the voxel grid center for the 3D metric computations. For the RealEstate10k dataset experiments, we used the NDC camera intrinsics which is provided, and for the NYU Depth V2 dataset experiments, we used the NDC camera intrinsics that are obtained for each scene, derived from the reconstruction process. In our 2.5D study, ground truth representations and DFM-aggregated point clouds (with associated confidences) were culled within the camera’s frustum determined by the camera’s pose and intrinsic parameters. Occluded points were removed using Z-culling with a depth buffer, ensuring only the closest points to the camera’s projection were retained  [^1] https://developer.nvidia.com/gpugems/gpugems/part-v-performance-and-practicalities/chapter-29-efficient-occlusion-culling. These points were scaled and voxelized into a grid of size $10\times 10\times 10$ from the 3D grid. The voxel grid was indexed with the input camera positioned at a 5-unit offset from the grid center along the $+z$-axis  [^2] https://docs.opencv.org/2.4/modules/calib3d/doc/camera_calibration_and_3d_reconstruction.html. The projected pixels and the associated confidences were used for 2D analysis, while the corresponding 3D points were incorporated into the 2.5D study for voxelization. Pixels without any projections are replaced with zero confidence.
@@
At the time of conducting these experiments, we used the publicly available Stable Diffusion XL Inpainting pipeline from the Hugging Face Diffusers library [75] with the RunwayML Stable Diffusion inpainting model. Note that future changes to this model or updates may affect reproducibility. During the 2D generation process with SDXL, we used a fixed set of seeds for all experiments. The input image is placed at the center of a $360\times 640$ canvas with 140 pixels of masking on each side of the center $360\times 360$ crop, by the image dimension of $360\times 140$. An additional filter was applied to remove cases where person(s) were detected with a confidence score of $\geq 0.5$. Both the image and mask are resized to $512\times 512$, and was resized back to $360\times 640$, to match the target dimension. In certain cases, the combination of seed and prompt led to the generation of NSFW content, which the model automatically rejected, returning a blank image for the outpainted regions. These samples were excluded from our analysis. To ensure fairness in aggregation, we resampled till we hit the target number for all sets in every prompt regime. We used 10 different prompts in three different regimes, (1) with object cues in text prompts, (2) without object cues in text prompts and (3) no text prompts. The prompts used are tabulated in:
@@
Parentheses indicate object cues, which are included in prompts with cues and omitted in prompts without cues. In the no-text-prompt regime, the prompt field is left empty. To ensure consistency in sampling for the ablation study, we generate 200 such valid cases.

### v3

@@ 3 Unobserved Object Detection
The task of unobserved object detection, as shown in Fig. 1, is to detect objects that are in the scene, but are not captured within the camera frustum. The approaches considered in this paper address this task by predicting a conditional distribution $\mathcal{D}_{\mathcal{I}}$ over a bounded space and set of semantic labels given a single RGB image $\mathcal{I}$, referred to as a spatio-semantic distribution (SSD). This facilitates the inference of the probable locations of unobserved objects outside the visible region captured by the image and its associated camera frustum, using contextual cues from the image. The spatial domain $\mathcal{X}$ defines the support of this distribution. The SSD is an estimate of the ground-truth conditional distribution $\Gamma_{\mathcal{I}}$. While this distribution is not observable, we have access to additional posed images from a realization of this conditional distribution for the purpose of evaluation. For each object label $o$, we denote the associated SSD as $\mathcal{D}_{\mathcal{I}o}$. This maps each $x\in\mathcal{X}$ to a likelihood in $[0,1]$. That is, the likelihood of detecting object $o$ at location $x$. For example, it assigns a likelihood to the presence of a chair one meter in front of and two meters to the right of the camera, which may not be visible in the input image. Note that we consider all distributions to be normalized. We consider distributions over two discretized spatial domains $\mathcal{X}$: (1) a 2D domain $\mathbb{I}$, discretized as a pixel grid, corresponding to an expansion of the input image; and (2) a 3D domain $\mathbb{V}$, discretized as a voxel grid, extending beyond the input camera’s frustum. The 2D domain $\mathbb{I}$ symmetrically extends the $H\times W$ pixel grid of the input image to $H\times W^{\prime}$, where $W^{\prime}>W$. The 3D domain $\mathbb{V}$ expands the frustum to an expanded voxel grid aligned with the input camera pose.
@@ 5.1 Metrics for Unobserved Object Detection
for input image $\mathcal{I}$ and object $o$. The normalization factor $\log|\mathcal{X}|$ makes the metric independent of the grid resolution and ensures that a uniform distribution evaluates to unit entropy. This asymmetric measure penalizes the prediction of low object likelihood where the ground-truth indicates the presence of an object. However, as required, it does not penalize the prediction of high object likelihoods where the ground-truth has not seen an object. For the remaining metrics, we consider high-likelihood detections above a threshold $\tau$. The thresholded ground truth is denoted by $\mathcal{G}_{\mathcal{I}o}^{\tau}$ for all $x\in\mathcal{X}$ where $\mathcal{G}_{\mathcal{I}o}(x)>\tau$. Let the detection label $y_{\mathcal{I}o}$ be $1$ if $\mathcal{G}_{\mathcal{I}o}^{\tau}$ is non-empty and 00 otherwise. The thresholded prediction $\mathcal{D}_{\mathcal{I}o}^{\tau}$ and label $\hat{y}_{\mathcal{I}o}$ are defined over $\mathcal{D}_{\mathcal{I}o}$.
@@ 5.2 Experiment Setup
All models were evaluated on 10 indoor scenes from the RealEstate10k test set [87] and the NYU Depth V2 dataset [65], automatically selected based on object detection frequency. After filtering, each scene has an average of 190 images in RealEstate10k ($720\times 1280$, resized to $360\times 640$ for $\mathbb{I}$, and center-cropped to $360\times 360$ for $\mathcal{I}$) and 1359 images in NYU Depth V2 ($480\times 640$, center-cropped to $360\times 640$ for $\mathbb{I}$ and center-cropped to $360\times 360$ for $\mathcal{I}$). One randomly selected frame per scene was used as the input image $\mathcal{I}$ for evaluation. COLMAP [61] and GLOMAP [48] were used for camera pose estimation (for NYU Depth V2) and for reconstruction to obtain depth maps. Object annotations for each expanded $360\times 640$ image $\mathbb{I}$ were generated using YOLOv8x [53, 26], pre-trained on the MS-COCO dataset [34], retaining those with confidence greater than $\tau_{\text{conf}}$. The 2D ground-truth spatial distributions $\mathcal{G}_{\mathcal{I}o}^{\text{2D}}$ were computed by a softmax normalization of the detected object confidence logits. The 3D ground-truth spatial distributions $\mathcal{G}_{\mathcal{I}o}^{\text{3D}}$ were computed by back-projecting the object confidence logits using the 3D reconstructions, voxelizing with average pooling, then applying softmax normalization. The voxel grid has resolution $20\times 20\times 20$ with grid size $1.0$. The 2.5D ground truth spatial distributions $\mathcal{G}_{\mathcal{I}o}^{\text{2.5D}}$ were computed similarly, by back-projecting the confidence logits within the input camera frustum and using a voxel grid of resolution $10\times 10\times 10$, with a grid size of $1.0$. Note that we only retain non-empty ground-truth distributions in the dataset: if a scene $s$ does not contain an object $o$, $(s,o)$ is not part of the dataset. See Appendices A and B for details on automatic scene pre-processing, filtering, selection, and reconstruction.
@@
The hyperparameters are set as follows: the classification threshold $\tau=1.4/|\mathcal{X}|$, the minimum detection confidence $\tau_{\text{conf}}=0.1$, and the number of subdomains $\mathcal{X}^{i}$ for region-wise accuracy $\ell=3$, corresponding to the $360\times 360$ input image $\mathcal{I}$, the $360\times 140$ domain to the left of the image and the $360\times 140$ domain to the right of the image. For DFM [73] inference, we use autoregressive sampling with 3 timesteps per frame, generating 50 intermediate frames with default settings: a temperature of $0.85$ and a guidance scale of $2.0$. Generating each sample took $25$ minutes per target pose per sample on average; that is, 48 days to generate all samples for each dataset.
For Stable Diffusion sampling [51], the Diffusers library SDXL inpainting model was employed. Input images were padded by 140 pixels on the left and right sides for masked inpainting. The temperature was set to $0.25$, the prompt guidance scales were drawn from $\{0.25,0.5,0.75,1.0\}$, the seeds from $\{0,42,123,999,2021\}$, and the prompts from a set of $10$, yielding $4\times 5\times 10=200$ samples per scene. For the VLMs, we used ChatGPT-4o [46], Claude-Sonnet-3.5 [3], and Gemini 1.5 Ultra [19] via their official APIs, and LLaVa-v1.6-34b [37, 38] via the Replicate API. The temperature was set to 0.25, the top-p value was set to 1.0 where applicable, and the order of the image–question pairs was randomly shuffled. The VLMs were queried 100 times per object, region, and scene, resulting in 6000 calls each ($20\,\text{images}\times 30\,\text{questions}\times 10\,\text{shuffles}$). Data acquisition and post-processing through these APIs took $\sim 8$h. Details are in Appendices C and D. Future updates or access restrictions to APIs used in this study could affect reproducibility. All experiments were run on one NVIDIA RTX A6000 GPU. Our code and evaluation protocols are on the project page [^1] https://1ssb.github.io/UOD/.
@@ 5.3 Results
Uniform 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1.000$ 0.553  $\pm$0.220 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1.000$ 0.558  $\pm$0.214
@@ 2.5D Task 3D Task
Uniform 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1.000$ 1.000  $\pm$0.000 1.000  $\pm$0.000 0.000  $\pm$0.000 $1.000$ 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1.000$ 1.000  $\pm$0.000 1.000  $\pm$$0$infty $1.000$
@@
Qualitative results are shown in Fig. 5(d), with the $360\times 360$ center-cropped input image $\mathcal{I}$ shown within the dashed lines in the first column, a top-down view of the DFM-predicted 3D distribution in the second column, the SDXL-predicted 2D distribution in the third column, and the ChatGPT-predicted 2D distribution in the last column. In particular, the DFM-based model usually predicts high likelihoods near the ground-truth object’s 3D location, including when the object is occluded or out-of-frame, which is a significantly harder task than the 2D prediction task; see, for example, the final row. More qualitative results can be found in Appendix E. Failure cases are shown in Fig. 6, where all models perform poorly and an object category (door) that elicits many false negatives in the object detector. We also observe that the diffusion-based outpainting approach has its own modes of failure, also shown in Fig. 6, where the results are not informative in a scene if the model is not prompted with the label of the object to be detected.
@@ 6 Conclusion
This work has several significant limitations. In particular, the inference time of 3D diffusion models, is currently very large and are costly to re-train. Since our approaches require many samples to approximate the conditional distribution, this is a significant bottleneck and stymies their use in real-time applications like in robotics. Another limitation is the need for object prompts in outpainting-based methods, since this is likely to bias the obtained empirical conditional distribution. Also, VLM-based evaluations may not be fully reliable, since it is possible these models could have seen and memorized the test images during training. Since the task requires that the model not over-fit to the ground truth realization—which is an instance of the true conditional distribution—there are inherent challenges in measuring performance that reflects accurate modeling of the underlying true distribution that is unavailable in real-world settings.
Future applications will apply the proposed pipeline to outdoor and large-scale scenes [36]. Synthetic datasets [12] can let us access known conditional distributions in simulation [15]. Beyond object detection, different spatio-semantic features like occupancy or affordances [23] are promising next steps. Improvements in solving this task could pave the way for robotics applications in reactive control [47], visual search [4], planning [27], navigation [17], and exploration [24]. This motivates the current work as a significant step toward improving machine vision capabilities for unobserved object detection.
Acknowledgements
Subhransu is supported by the University Research Scholarship at the Australian National University. This work is also supported in part by the DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) program under grant agreement number HR00112490421.
@@ Appendix A Object, Scene and Parameter Selection Methodology
To systematically select diverse scenes, we sampled video frames in the RealEstate10k test dataset, which contains world-to-camera poses and pinhole camera intrinsics [87], built using the ORBSLAM pipeline (with an ambiguous scale) [42, 43]. Due to memory limitations, we downloaded 250 random scenes for analysis. RealEstate10k contains indoor and outdoor scenes but lacks metadata or labels, thus to interpret and filter each scene, our objective was to select representative viewpoints in each scene. We selected these frames, for every scene, at approximately equidistant [9] intervals in $\mathrm{SE}(3)$, between the two farthest camera-to-world poses for each scene. Using these selected images, per scene, we used the ChatGPT-4o API [46] to classify scenes as indoor or outdoor based on these 10 images per scene. For the indoor scenes, we further utilized the API to categorize them into four types: kitchen, study, bedroom, and living room. This automated labeling ensured that our subsequent analysis remained focused and consistent across the relevant indoor scenes.
@@
We deployed an end-to-end automated object selection pipeline on the randomly acquired scenes, logging the occurrence frequency of detected objects across the subsets of both datasets. From the combined scenes of NYU Depth V2 and RealEstate10k scenes, we selected the 20 most frequently detected object classes among the 80 classes in the COCO dataset [34]. For our analysis, we focused on 10 random object categories from these 20 classes that were present in at least one of the chosen scenes in the combined choices of scenes from both datasets. The COCO object categories analyzed include: refrigerator, TV, bed, chair, sink, oven, book, laptop, couch, and door. We used Yolov8x [53, 26] for its strong adaptability for images of various resolutions and superior generalizability. Note that all 2D/3D cases analyzed were manually inspected and verified against YOLOv8x annotations to ensure high quality ground truth annotations.
@@ Appendix C Additional implementation details
The experiments were conducted on a single machine equipped with an NVIDIA A6000 GPU, an AMD Ryzen Threadripper Pro 5995WX CPU, and 64GB of DDR5 RAM, running Ubuntu 22.04 LTS. From sampling to the final metrics summary, our end-to-end pipeline takes up to 38 minutes utilizing multiprocessing.
@@
For the 3D study, we utilized the aggregated point cloud from the samples from the corresponding poses, and correspondingly all valid 3D points in the ground-truth reconstruction, including occluded parts of the scene, voxelized into a grid of size $20\times 20\times 20$, such that the camera position is at the center of the grid. For the SDXL-based study, taking a similar approach as Chen et al. [11] to lift our representation to 2.5D, we generated metric depths up to the far plane of $10$ units using DepthAnythingv2 [80, 81], selected for its superior cross-domain generalizability and compatibility with our known camera model. This approach provided depths within the frustum, ensuring scale consistency across the analyses, making sure that the camera centers are aligned to the voxel grid center for the 3D metric computations. For the RealEstate10k dataset experiments, we used the NDC camera intrinsics which is provided, and for the NYU Depth V2 dataset experiments, we used the NDC camera intrinsics that are obtained for each scene, derived from the reconstruction process. In our 2.5D study, ground truth representations and DFM-aggregated point clouds (with associated confidences) were culled within the camera’s frustum determined by the camera’s pose and intrinsic parameters. Occluded points were removed using Z-culling with a depth buffer, ensuring only the closest points to the camera’s projection were retained  [^2] Efficient occlusion culling methods. These points were scaled and voxelized into a grid of size $10\times 10\times 10$ from the 3D grid. The voxel grid was indexed with the input camera positioned at a 5-unit offset from the grid center along the $+z$-axis  [^3] Camera Calibration and 3D reconstruction documentation. The projected pixels and the associated confidences were used for 2D analysis, while the corresponding 3D points were incorporated into the 2.5D study for voxelization. Pixels without any projections are replaced with zero confidence.
@@
For our experiments, we used the publicly available Stable Diffusion XL Inpainting pipeline from the Hugging Face Diffusers library [75]. Future changes to APIs and access to this model may affect reproducibility. During the 2D generation process with SDXL, we used a fixed set of seeds for all experiments. The input image is placed at the center of a $360\times 640$ canvas with 140 pixels of masking on each side of the center $360\times 360$ crop, by the image dimension of $360\times 140$. An additional filter was applied to remove cases where person(s) were detected with a confidence score of $\geq 0.5$. Both the image and mask are resized to $512\times 512$, and was resized back to $360\times 640$, to match the target dimension. In certain cases, the combination of seed and prompt led to the generation of NSFW content, which the model automatically rejected, returning a blank image for the outpainted regions. These samples were excluded from our analysis. To ensure fairness in aggregation, we resampled till we hit the target number for all sets in every prompt regime. We used 10 different prompts in three different regimes, (1) with object cues in text prompts, (2) without object cues in text prompts and (3) no text prompts. The prompts used were:
@@
Parentheses indicate object cues, which are included in prompts with cues and omitted in prompts without cues. In the no-text-prompt regime, the prompt field is left empty. To ensure consistency in sampling for the ablation study, we generate 200 such prompts in total.
CAR
(a) $\mathcal{I}$ (center) and $\mathbb{I}$ (full)
(b) DFM-based $\mathcal{D}_{\mathcal{I}o}^{\text{3D}}$
(c) SDXL-based $\mathcal{D}_{\mathcal{I}o}^{\text{2D}}$
(d) ChatGPT-4o-based $\mathcal{D}_{\mathcal{I}o}^{\text{2D}}$
Figure 11: Results for an outdoor scene–object pair. The GT position is denoted by a white star.
0 [Uncaptioned image] 1